# Diving into the Fusion of Monocular Priors for Generalized Stereo Matching

### v1

@@ Abstract
The matching formulation makes it naturally hard for the stereo matching to handle ill-posed regions like occlusions and non-Lambertian surfaces. Fusing monocular priors has been proven helpful for ill-posed matching, but the biased monocular prior learned from small stereo datasets constrains the generalization. Recently, stereo matching has progressed by leveraging the unbiased monocular prior from the vision foundation model (VFM) to improve the generalization in ill-posed regions. We dive into the fusion process and observe three main problems limiting the fusion of the VFM monocular prior. The first problem is the misalignment between affine-invariant relative monocular depth and absolute depth of disparity. Besides, when we use the monocular feature in an iterative update structure, the over-confidence in the disparity update leads to local optima results. A direct fusion of a monocular depth map could alleviate the local optima problem, but noisy disparity results computed at the first several iterations will misguide the fusion. In this paper, we propose a binary local ordering map to guide the fusion, which converts the depth map into a binary relative format, unifying the relative and absolute depth representation. The computed local ordering map is also used to re-weight the initial disparity update, resolving the local optima and noisy problem. In addition, we formulate the final direct fusion of monocular depth to the disparity as a registration problem, where a pixel-wise linear regression module can globally and adaptively align them. Our method fully exploits the monocular prior to support stereo matching results effectively and efficiently. We significantly improve the performance from the experiments when generalizing from SceneFlow to Middlebury and Booster datasets while barely reducing the efficiency. https://github.com/YaoChengTang/Diving-into-the-Fusion-of-Monocular-Priors-for-Generalized-Stereo-Matching
@@ 1 Introduction
Taking advantage of large-scale scenes and the easily collected ground truth of monocular depth, the vision foundation model can provide an unbiased monocular prior [16, 48, 49]. Recently, some methods have made great progress in fusing the unbiased monocular prior into the stereo matching to improve the generalization in ill-posed regions [8, 3, 46]. In this paper, we dive into the fusion mechanism and find three main problems limiting a full exploration of the unbiased monocular prior. The first problem lies in the natural gap between the affine-invariant relative depth from monocular depth and absolute depth from disparity. Although we can forcibly align the two kinds of depth with a complex mutual refinement, these alignments could involve heavy computation and greatly harm the efficiency [8, 46]. The other problem exists in the fusion with monocular feature maps in an iterative refinement structure [47, 43, 7]. The implicit feature fusion makes the fusion more biased to the binocular information due to the iterative update training scheme, where the over-confidence of the disparity update causes local optima, as shown in Figure 1. An additional fusion of monocular depth could alleviate the local optima, but the direct fusion of the depth map is easily affected by the noisy depth results. Even with unbiased and smooth monocular depth from the VFM, noisy disparity at early iterations slows down a good fusion.
@@
We compare our model with state-of-the-art methods using the standard setting, training on the SceneFlow dataset, and testing on five real-world datasets with various ill-posed areas, including KITTI 2012, KITTI 2015, Middlebury, ETH3D, and Booster. The results demonstrate that our method significantly enhances the performance of state-of-the-art approaches, as shown in Figure 1. Experiments show that our method achieves a 10-point improvement in the bad2 metric for transparent regions on the Booster dataset and reduces errors by more than 50% on Middlebury and ETH3D, where we do not use additional stereo data or specific data augmentation. Meanwhile, even involving a VTF model, our method barely raises the time cost, benefiting from the elegant explicit designs.
@@ 2.1 Generalized Stereo Matching
Generalized stereo matching aims to produce a reliable dense disparity map when the target domain (e.g., real-world data) differs from the source domain (e.g., synthetic data). Some methods focus on the learning of domain invariant features [4, 22, 9, 54, 33, 17, 40]. MS-PSMNet [4] replaces the learning-based features with hand-crafted features to force the stereo network focus on the matching space. MS-PSMNet has achieved great improvement, but its hand-crafted features limit the performance of the stereo network. Thus, many methods turn to improve the training process by transforming learning [22], meta-learning [40, 17], contrastive learning [54, 33], and fisher information [9].
@@
The above feature-based methods significantly improve the generalization of stereo matching. However, it is difficult for them to eliminate domain gaps due to the complexity of real-world scenarios. Then, some researchers integrate other modalities to enrich the features of RGB images [52, 42]. They achieve impressive performance but require additional devices. Instead, other researchers propose to generate more and better data for training [39, 5, 41, 3, 46]. AdaStereo [39] and HVT-RAFT [5] augment the training data in color space to enrich the domain distribution in the synthetic dataset. They improve greatly, but the rendered images in the synthetic dataset are unrealistic to the real world. Thus, NerfStereo [41] turns to reconstruct the real-world scenes from Nerf and re-renders stereo images to improve the quality of training data.
In addition to improving generalization from features and data, some methods focus on architecture design with specific knowledge of stereo matching [53, 20, 14, 47, 7, 43, 12]. DSMNet [53] uses long-range matching information in cost aggregation to correct the mismatched points. The improvement in cost aggregation is remarkable, but the additional operations in 3D space are time-consuming. Many methods then turn to incorporating global information when constructing cost volume. STTR [20] and CSTR [14] use transformers to capture long-range matching information. Other methods [43, 47, 7] build auxiliary volume to augment the original cost volume. There are also some methods resolving the generalization problem from uncertainty learning [15, 23].
@@ 2.2 Fusing Monocular and Stereo Estimation
Inspired by the human vision system that fuses binocular disparity and monocular depth cues [34, 10, 44, 45, 8, 3, 46], researchers are exploring the fusion mechanism for machine vision. Tradition method [35] achieves it via an MRF optimization objective that relies on the binocular disparity and monocular cues. Deep learning methods mainly focus on volume fusion or depth map fusion [8, 3, 46]. The volume fusion methods introduce monocular priors into cost volume [51, 19, 8, 3, 46]. These methods require a fixed disparity range and depend on domain-biased monocular priors. Our method assumes no limit on the disparity range and introduces unbiased monocular priors from a pre-trained large model. The depth map fusion methods fuse the monocular depth and the binocular disparity in one-stage post-processing [24, 6, 2, 55, 1]. These methods mainly predict affine-invariant monocular depth, which is not aligned with binocular disparity and results in noise in fusion. Instead, our method uses local ordering maps to make compatibility between monocular depth and binocular disparity, reducing noise in the iterative matching process. Based on the optimized binocular disparity, the monocular depth is further globally aligned with it by learning two parameters to solve the scale ambiguity in monocular results.
@@ 3.1 Monocular Encoder
As shown in Figure 2, given an image with a resolution of $H\times W$, we pre-process the image as DepthAnything v2 [49] by resizing the longest side of the image to 512 pixels. The resized image is then fed into a frozen DepthAnything v2 to extract intermediate features before the DPTHead [32] and monocular depth after the DPTHead. These intermediate features and monocular depth are subsequently resized to a $H/4\times W/4$ resolution using a bilinear function to interact with the stereo-matching pipeline. We build a two-stream convolution module to generate the initial hidden state and monocular context features from the intermediate features. Although the output of DepthAnything v2 is inverse depth (disparity under unknown camera parameters), we refer to it as monocular depth for consistency description.
@@ 3.2 Iterative Local Fusion
The iterative local fusion module leverages the binary local ordering map to update disparity with monocular priors iteratively. The binary local ordering map helps mitigate the impact of outlier noises by converting absolute values into ordering relationships, which is much more robust than the pixel-wise depth value. Besides, it also unifies the affine-invariant depth and absolute disparity to be compatible with the order relationship.
To compute the binary local ordering map, we use a series of LBP-like operations [29, 28] with varying window sizes to extract local ordering features. Each LBP-like operation consists of a convolution with fixed weights followed by a sigmoid function, which measures the relative depth relationships between the center pixel of the window and its neighboring pixels, indicating which pixels are closer or farther. The formulation of the local ordering map $M_{O}$ at pixel $(u,v)$ is as follows:
$M_{O}(u,v)=\{\sigma(D(u^{\prime},v^{\prime})-D(u,v))\ |\ (u^{\prime},v^{\prime})\in\mathcal{N}_{(u,v)}\},$ (1)
where $D$ is depth or disparity map, $\sigma$ is the sigmoid function and $\mathcal{N}$ is the neighborhood. To employ the binary local ordering map into the iterative refinement structure, we use the LBP-like encoder to extract local ordering maps from both monocular depth and binocular disparity in the previous iteration, as shown in Figure 2. These two kinds of local ordering maps are concatenated to predict the monocular guidance. The guidance $G$ is then used to re-weight the initial disparity update $\Delta_{d}$ to avoid local optima. We represent the monocular guidance as a Beta distribution $Beta(\alpha,\beta)$. During training, $G$ is sampled via reparameterization trick. At test time, we compute the distribution expectation to obtain it: $G=\alpha/(\alpha+\beta)$. $G$ is used to reweight the disparity update, as defined in Equation (1) of the main text.
@@ 3.3 Global Fusion
After all iterations of disparity update, we use a global fusion module to incorporate fine-grained 3D shape priors from the monocular depth map into the disparity map, as shown in Figure 2. Here, we formulate the optimized binocular disparity as a registered version of monocular depth with minor noise by specific intrinsic parameters. Therefore, monocular depth can be globally registered to binocular disparity. The registration can be deemed a linear regression problem with noise between monocular depth and binocular disparity. To this end, we first align the monocular depth $D_{m}$ with the optimized disparity $D_{d}$ by estimating two registration parameters, ${a,b}$ by
@@
$D_{f}=c\cdot D_{d}^{T}+(1-c)\cdot\tilde{D}_{m},$ (5)
where $c$ is a confidence map. $D_{f}$ is the final disparity prediction.
@@ 4.1 Implementation Details
For the stereo part, our pipeline is built on the classical iterative structure of RAFT-Stereo [21], which is widely used and flexible to deploy without stacking network tricks to raise the computation burden. As for the monocular part, we use DepthAnything V2 [48, 49] to extract unbiased monocular priors. Still, it is flexible to use other VTFs as long as the monocular prior is not overfitted. We set parameters set as $r=1$ and $\gamma=0.9$. The window sizes for the LBP-like operations are configured to ${5,3}$. The training was conducted on 4 NVIDIA A40 GPUs using the AdamW optimizer with a one-cycle learning rate schedule. During training, the DepthAnything V2 module remains frozen. Specifically, we first train the model without the global fusion module on the SceneFlow dataset, using a maximum learning rate of 0.0002, a batch size of 8, and for 100k steps, maintaining the consistency of matching parts with the total data used in RAFT-Stereo. Then, we train the monocular registration of the global fusion module while keeping the other modules frozen, using a maximum learning rate of 0.0005 and a batch size of 32 for 100k steps on the SceneFlow dataset. Finally, we train the entire global fusion module while keeping the other modules frozen, using a maximum learning rate of 0.0005, a batch size of 32, and 100k steps on the SceneFlow dataset. Our results are not sensible to the hyper-parameters of the training process. With the training and testing codes provided in the supplementary materials, all the evaluation results can simply be reproduced.
Method Year Additional Data/Aug KITTI 2015 KITTI 2012 Middlebury (H) ETH3D
@@
Table 1: Generalization from SceneFlow dataset to KITTI2015, KITTI 2012, Middlebury (H), and ETH3D dataset. ’ME’ represents our monocular encoder module. * represents the results evaluated in our metrics and settings using official models and weights. ’All’, ’NonOcc’, and ’Occ’ represent all regions, non-occluded regions, and occluded regions, respectively.
Method Additional Data/Aug Booster (Q)
@@
Table 2: Generalization from SceneFlow dataset to Booster dataset in quarter resolution and balanced set. ’ME’ represents our monocular encoder module. ’All’, ’Trans’, and ’NonTrans’ represent all regions, transparent regions, and nontransparent regions, respectively.
Method Additional Data/Aug DrivingStereo
@@
RAFTStereo 1.01 0.97 1.8 0.95
IGEV 1.11 1.11 2.32 1.14
MoCha-Stereo 1.01 0.99 1.35 0.98
Selective-IGEV 1.18 1.13 2.22 1.12
NerfStereo 0.90 0.91 1.46 1.01
@@
$\delta^{1}$uparrow RMS $\downarrow$ $\delta^{1}$uparrow RMS $\downarrow$ $\delta^{1}$uparrow RMS $\downarrow$
@@
Table 4: Comparison to DepthAnything V2 and Metric3D on Middlebury. ’M’: the fine-tuned metric version. ’GA’: alignment using the same registration parameters from GT for all pixels.
4.2 Evaluation
Metrics. (1) We use two metrics: EPE, which measures the mean absolute disparity error in pixels, and Bad $x$, which represents the percentage of pixels where the predicted disparity deviates from the ground truth by at least $x$ pixels. (2) It is important to note that many recent methods report their results with some implicit assumptions, such as evaluating only pixels with ground truth disparity less than 192 or only evaluating non-occluded regions. In our experiments, unless otherwise specified, both ours and compared methods consider all regions as the classical metric does without limitations. For the Middlebury dataset, we evaluate both all regions and non-occluded regions. For the Booster dataset, we evaluate all regions, as well as transparent and non-transparent regions. (3) Additionally, we observe fluctuations in model performance when trained with different numbers of steps. To fully analyze the improvement contributed by each model component, we calculate the mean and standard deviation (std) of results from the last 100k, 90k, and 80k training steps. We use $mean\pm std$ to measure the accuracy and robustness of our model.
Exp Middlebury
@@ 4.2 Evaluation
Baseline 2.11$\pm$0.16 14.12$\pm$0.64
Baseline w/o mono feature 1.83$\pm$0.11 12.45$\pm$0.86
Baseline + ME 1.42$\pm$0.01 9.81$\pm$0.18
Baseline + ME + IDF 1.41$\pm$0.04 10.34$\pm$0.19
Baseline + ME + PF 1.41$\pm$0.00 9.71$\pm$0.00
Baseline + ME + ILF 1.20$\pm$0.08 9.06$\pm$0.70
Baseline + ME + ILF + GF 1.15$\pm$0.01 8.35$\pm$0.04
Table 5: The effectiveness of each module. The baseline is RAFTStereo, while ME is our monocular encoder, DF is iterative direct fusion, ILF is iterative local fusion, PF is post-fusion, and GF is global fusion. w/o mono feature means removing the context features from RAFT-Stereo.
@@
4.3 In-the-wild Generalization Ability
Exp LBP Kernel S OP r Middlebury
@@ 4.3 In-the-wild Generalization Ability
L(1) 1 L 1 1.38$\pm$0.06 10.20$\pm$0.53
L(2) 3 L 1 1.36$\pm$0.04 10.10$\pm$0.25
L(3) 3 L 1 1.44$\pm$0.06 9.65$\pm$0.26
L(4) 5,3 L 1 1.20$\pm$0.08 9.06$\pm$0.70
L(5) 9,7,5,3 L 1 1.32$\pm$0.07 9.57$\pm$0.62
L(6) C 1 1.32$\pm$0.14 9.53$\pm$1.03
L(7) DC 1 1.39$\pm$0.03 9.71$\pm$0.29
L(8) 13,11,9,7,5,3 L 1 1.31$\pm$0.02 9.89$\pm$0.04
L(9) 5,3, L 2 1.32$\pm$0.03 9.45$\pm$0.14
L(10) 5,3 L 3 1.26$\pm$0.09 9.42$\pm$0.42
Table 6: Ablation study on iterative local fusion. ’S’ is sigmoid function in LBP-like operation. ’OP’ represents the type operation, ’L’ is the LBP-like operation, ’C’ is the convolution, ’DC’ is the deeper convolution, and ’r’ is the amplitude parameter.
Exp Reg Confidence Middlebury
@@
G(1) Cost 1.23$\pm$0.03 9.69$\pm$0.38
MonoDepth 1.19$\pm$0.02 8.72$\pm$0.20
G(2) Cost 1.18$\pm$0.02 8.77$\pm$0.07
G(3) Hybrid 1.15$\pm$0.01 8.35$\pm$0.04
Table 7: Ablation study on the global fusion. ’Reg’ means registration for monocular depth. ’Cost’ means we estimate the confidence from the sampled cost volume. ’Hybrid’ means we estimate the confidence from the concatenation of sampled cost volume, hidden state, and guidance from the last iteration. ’MonoDepth’ means evaluation of the registered monocular depth.

### v2

@@ Abstract
The matching formulation makes it naturally hard for the stereo matching to handle ill-posed regions like occlusions and non-Lambertian surfaces. Fusing monocular priors has been proven helpful for ill-posed matching, but the biased monocular prior learned from small stereo datasets constrains the generalization. Recently, stereo matching has progressed by leveraging the unbiased monocular prior from the vision foundation model (VFM) to improve the generalization in ill-posed regions. We dive into the fusion process and observe three main problems limiting the fusion of the VFM monocular prior. The first problem is the misalignment between affine-invariant relative monocular depth and absolute depth of disparity. Besides, when we use the monocular feature in an iterative update structure, the over-confidence in the disparity update leads to local optima results. A direct fusion of a monocular depth map could alleviate the local optima problem, but noisy disparity results computed at the first several iterations will misguide the fusion. In this paper, we propose a binary local ordering map to guide the fusion, which converts the depth map into a binary relative format, unifying the relative and absolute depth representation. The computed local ordering map is also used to re-weight the initial disparity update, resolving the local optima and noisy problem. In addition, we formulate the final direct fusion of monocular depth to the disparity as a registration problem, where a pixel-wise linear regression module can globally and adaptively align them. Our method fully exploits the monocular prior to support stereo matching results effectively and efficiently. We significantly improve the performance from the experiments when generalizing from SceneFlow to Middlebury and Booster datasets while barely reducing the efficiency.
[Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image]
@@ 1 Introduction
Taking advantage of large-scale scenes and the easily collected ground truth of monocular depth, the vision foundation model can provide an unbiased monocular prior [16, 52, 53]. Recently, some methods have made great progress in fusing the unbiased monocular prior into the stereo matching to improve the generalization in ill-posed regions [8, 3, 50]. In this paper, we dive into the fusion mechanism and find three main problems limiting a full exploration of the unbiased monocular prior. The first problem lies in the natural gap between the affine-invariant relative depth from monocular depth and absolute depth from disparity. Although we can forcibly align the two kinds of depth with a complex mutual refinement, these alignments could involve heavy computation and greatly harm the efficiency [8, 50]. The other problem exists in the fusion with monocular feature maps in an iterative refinement structure [51, 46, 7]. The implicit feature fusion makes the fusion more biased to the binocular information due to the iterative update training scheme, where the over-confidence of the disparity update causes local optima, as shown in Figure 1. An additional fusion of monocular depth could alleviate local optima, but direct fusion of the depth is easily affected by noisy depth results. Even with unbiased, smooth monocular depths from the VFM, noisy disparity in the first several iterations slows down good fusion.
@@
We compare our model with SOTA methods under the standard setting: training on SceneFlow and testing on five real-world datasets (KITTI 2012&2015, Middlebury, ETH3D, and Booster) with various ill-posed regions. Results demonstrate our method significantly boosts SOTA performance, as shown in Figure 1. Our method achieves a 10-point improvement in the bad2 metric for transparent regions on Booster and reduces errors by over 50% on Middlebury and ETH3D, without using extra stereo data or specific augmentation. Despite involving a VTF model, it barely increases the time cost due to the elegant designs.
@@ 2.1 Generalized Stereo Matching
Generalized stereo matching aims to produce a reliable dense disparity map when the target domain (e.g., real-world data) differs from the source domain (e.g., synthetic data). Some methods focus on learning domain-invariant features [4, 23, 9, 58, 35, 17, 42]. MS-PSMNet [4] replaces learning-based features with hand-crafted ones to force the stereo network to focus on the matching space. Although it achieves significant improvement, its hand-crafted features limit the performance. Thus, many methods improve training via transfer learning [23], meta-learning [42, 17], contrastive learning [58, 35], and Fisher information [9].
@@
The above feature-based methods significantly improve stereo matching generalization. However, due to real-world complexity, they struggle to eliminate domain gaps. Some researchers integrate other modalities to enrich RGB features [56, 44], achieving strong performance but requiring extra devices. Others instead generate more and better training data [41, 5, 43, 3, 50]. AdaStereo [41] and HVT-RAFT [5] augment data in color space to enrich domain distribution. While effective, rendered images remain unrealistic. Thus, NerfStereo [43] reconstructs real scenes via NeRF and re-renders stereo images to improve training quality.
Beyond improving generalization through features and data, some methods focus on architecture design leveraging stereo-specific knowledge [57, 21, 14, 51, 7, 46, 12]. DSMNet [57] uses long-range matching in cost aggregation to correct mismatches. While effective, its 3D operations are time-consuming. Many later methods incorporate global information in cost volume construction. STTR [21] and CSTR [14] apply transformers to capture long-range dependencies. Others [46, 51, 7] build auxiliary volumes to enhance the original cost volume. Some approaches also improve generalization via uncertainty learning [15, 25].
@@ 2.2 Fusing Monocular and Stereo Estimation
Inspired by the human visual system fusing binocular disparity and monocular cues [36, 10, 48, 49, 8, 3, 50], researchers have explored similar fusion mechanisms in machine vision. Traditional methods [37] use MRF optimization based on disparity and monocular cues. Deep learning methods mainly adopt volume or depth map fusion [8, 3, 50]. Volume fusion injects monocular priors into cost volumes [55, 20, 8, 3, 50], but relies on fixed disparity ranges and domain-biased priors. In contrast, our method removes disparity range limits and uses unbiased priors from a large pre-trained model. Depth map fusion methods [26, 6, 2, 59, 1] combine monocular and binocular results in post-processing, but often suffer from misalignment and noise due to affine-invariant monocular predictions. Instead, we employ local ordering maps for better compatibility, reducing noise during matching. The final monocular depth is globally aligned to the optimized disparity by learning two parameters to address scale ambiguity.
@@ 3.1 Monocular Encoder
As shown in Figure 2, given an image with a resolution of $H\times W$, we pre-process the image as DepthAnything v2 [53] by resizing the longest side of the image to 512 pixels. The resized image is then fed into a frozen DepthAnything v2 to extract intermediate features before the DPTHead [34] and monocular depth after the DPTHead. These intermediate features and monocular depth are subsequently resized to a $H/4\times W/4$ resolution using a bilinear function to interact with the stereo-matching pipeline. We build a two-stream convolution module to generate the initial hidden state and monocular context features from the intermediate features. Although both the prediction and supervision of DepthAnything v2 are in the form of inverse depth, which is disparity under some unknown camera parameters and baseline, we still refer to it as monocular depth here to maintain consistency with the terminology used in DepthAnything v2.
@@ 3.2 Iterative Local Fusion
The iterative local fusion module leverages the binary local ordering map to update disparity with monocular priors iteratively. The binary ordering map $M_{O}$ encodes the relative ordering of depth $D$ between a center pixel and its neighbors:
$M_{O}(u,v)=\{\sigma(D(u^{\prime},v^{\prime})-D(u,v))\},$ (1)
where $(u^{\prime},v^{\prime})\in\mathcal{N}_{(u,v)}$, $\sigma$ is the sigmoid function and $\mathcal{N}$ is the neighborhood. The binary local ordering map helps mitigate the impact of outlier noises by converting absolute values into ordering relationships, which is much more robust than the pixel-wise depth value. Besides, it also unifies the affine-invariant depth and absolute disparity to be compatible with the order relationship.
To compute the binary local ordering map, we use a series of LBP-like operations [31, 30] with varying window sizes to extract local ordering features. Each LBP-like operation consists of a convolution with fixed weights followed by a sigmoid function, which measures the relative depth relationships between the center pixel of the window and its neighboring pixels, indicating which pixels are closer or farther. To employ the binary local ordering map into the iterative refinement structure, we use the LBP-like encoder to extract local ordering maps from both monocular depth and binocular disparity in the previous iteration, as shown in Figure 2. These two kinds of local ordering maps are concatenated to predict the monocular guidance. The guidance $G$ is modeled by a Beta distribution with parameters $\{\alpha,\beta\}$ predicted via convolutions. During training, $G$ is sampled via reparameterization: $G=g_{1}/(g_{1}+g_{2})$, $g_{1}\sim\text{Gamma}(\alpha,1)$, $g_{2}\sim\text{Gamma}(\beta,1)$. At test time, $G=\alpha/(\alpha+\beta)$. The guidance $G$ is then used to re-weight the initial disparity update $\Delta_{d}$ to avoid local optima.
@@ 3.3 Global Fusion
After all iterations of disparity update, we use a global fusion module to incorporate fine-grained 3D shape priors from the monocular depth map into the disparity map, as shown in Figure 2. Here, we formulate the optimized binocular disparity as a registered version of monocular depth with minor noise by specific intrinsic parameters. Therefore, monocular depth can be globally registered to binocular disparity. The registration can be deemed a linear regression problem with noise between monocular depth and binocular disparity. To this end, we first align the monocular depth $D_{m}$ with the optimized disparity $D_{d}$ by estimating two registration parameters, a,b{a,b} by
@@
$D_{f}=c\cdot D_{d}^{T}+(1-c)\cdot\tilde{D}_{m}.$ (5)
$D_{f}$ is the final disparity prediction.
@@ 4.1 Implementation Details
For the stereo part, our pipeline is built on the classical iterative structure of RAFT-Stereo [22], which is widely used and flexible to deploy without stacking network tricks to raise the computation burden. As for the monocular part, we use DepthAnything V2 [52, 53] to extract unbiased monocular priors. Still, it is flexible to use other VTFs as long as the monocular prior is generalizable to practical scenarios. We set parameters as $r=1$ and $\gamma=0.9$. The window sizes for the LBP-like operations are configured to 5,3{5,3}. The training was conducted on 4 NVIDIA A40 GPUs using the AdamW optimizer with a one-cycle learning rate schedule. During training, the DepthAnything V2 module remains frozen. Specifically, we first train the model without the global fusion module on the SceneFlow dataset, using a maximum learning rate of 0.0002, a batch size of 8, and for 100k steps, maintaining the consistency of matching parts with the total data used in RAFT-Stereo. Then, we train the monocular registration of the global fusion module while keeping the other modules frozen, using a maximum learning rate of 0.0005 and a batch size of 32 for 100k steps on the SceneFlow dataset. Finally, we train the entire global fusion module while keeping the other modules frozen, using a maximum learning rate of 0.0005, a batch size of 32, and 100k steps on the SceneFlow dataset. Our results are not sensitive to the hyperparameters of the training process.
Method Year
Additional
Data/Aug
KITTI 2015 KITTI 2012 Middlebury (H) ETH3D
@@
Table 1: Generalization from SceneFlow dataset to KITTI2015, KITTI 2012, Middlebury (H), and ETH3D dataset. ‘ME’ represents our monocular encoder module. * represents the results evaluated in our metrics and settings using official models and weights. ‘All’, ‘NonOcc’, and ‘Occ’ represent all regions, non-occluded regions, and occluded regions, respectively.
Method
Additional
Data/Aug
Booster (Q)
@@
Table 2: Generalization from SceneFlow dataset to Booster dataset in quarter resolution and balanced set. ‘ME’ represents our monocular encoder module. ‘All’, ‘Trans’, and ‘NonTrans’ represent all regions, transparent regions, and nontransparent regions, respectively.
4.2 Evaluation
Metrics. (1) We use two metrics: EPE, which measures the mean absolute disparity error in pixels, and Bad $x$, which represents the percentage of pixels where the predicted disparity deviates from the ground truth by at least $x$ pixels. (2) It is important to note that many recent methods report their results with some implicit assumptions, such as evaluating only pixels with ground truth disparity less than 192 or only evaluating non-occluded regions. In our experiments, unless otherwise specified, both ours and the compared methods consider all regions as the classical metric does without limitations. For the Middlebury dataset, we evaluate both all regions and non-occluded regions. For the Booster dataset, we evaluate all regions, as well as transparent and non-transparent regions. (3) Additionally, we observe fluctuations in model performance when trained with different numbers of steps. To fully analyze the improvement contributed by each model component, we calculate the mean and standard deviation (std) of results from the last 100k, 90k, and 80k training steps. We use $mean\pm std$ to measure the accuracy and robustness of our model.
4.3 In-the-wild Generalization Ability
Method
Additional
Data/Aug
DrivingStereo
@@ 4.3 In-the-wild Generalization Ability
RAFTStereo [22] 1.01 0.97 1.8 0.95
IGEV [51] 1.11 1.11 2.32 1.14
MoCha-Stereo [7] 1.01 0.99 1.35 0.98
Selective-IGEV [46] 1.18 1.13 2.22 1.12
NerfStereo [43] 0.90 0.91 1.46 1.01
@@
$\delta^{1}$ $\uparrow$ RMS $\downarrow$ $\delta^{1}$ $\uparrow$ RMS $\downarrow$ $\delta^{1}$ $\uparrow$ RMS $\downarrow$
@@
Table 4: Comparison to DepthAnything V2 and Metric3D on Middlebury (H). M: the fine-tuned metric version. GA: alignment to GT scale using identical registration parameters computed from GT for all pixels.
Exp Middlebury (H)
@@
[…]
Table 5: The effectiveness of each module. The baseline is RAFTStereo. w/o mono feature: removing the context features from RAFTStereo. ME: our monocular encoder, DF: iterative direct fusion, ILF: iterative local fusion, PF: post-fusion, GF: global fusion.
@@
Exp
LBP
Kernel
S OP r Middlebury (H)
@@
[…]
Table 6: Ablation study on iterative local fusion. S: an LBP-like operation with or without a sigmoid function. OP: the type of operation, L: the LBP-like operation, C: the convolution, DC: a deeper convolution, r: the amplitude parameter.
Exp Reg Confidence Middlebury (H)
@@
[…]
Table 7: Ablation study on the global fusion. Reg: registration for monocular depth. Cost: estimating the confidence from the sampled cost volume. Hybrid: estimating the confidence from the concatenation of sampled cost volume, hidden state, and guidance from the last iteration. MonoDepth: evaluation of the registered monocular depth.